OpenStudy (daisiablis11):

I'm confused as to how 0 - ln(sqrt(2)/2)) = ln(2/sqrt(2)) = ln(sqrt(2))?

satellite73 (satellite73):

which equality is unclear?

OpenStudy (daisiablis11):

The whole thing.

satellite73 (satellite73):

ok lets go slow

satellite73 (satellite73):

first of all the zero is unnecessary right? \[0-\ln(\frac{\sqrt2}{2})=-\ln(\frac{\sqrt2}{2})\]

satellite73 (satellite73):

now is it clear that \[-\ln(x)=\ln(\frac{1}{x})\]?

OpenStudy (daisiablis11):

yes.

satellite73 (satellite73):

there are lots of explanations if it is not obvious to you one is that \[\ln(x^n)=n\ln(x)\] so \[-1\ln(x)=\ln(x^{-1})=\ln(\frac{1}{x})\] but there are other ways to see it

satellite73 (satellite73):

ok so if \(x=\frac{\sqrt2}{2}\) then \[\frac{1}{x}=\frac{2}{\sqrt2}\]that should make the first equality clear

OpenStudy (daisiablis11):

Yes. I understand to that part now.

satellite73 (satellite73):

ok then for the next equality, it is always true that \[\frac{a}{\sqrt a}=\sqrt{a}\]

satellite73 (satellite73):

there are lots of ways to see that too one is two square \[\frac{a}{\sqrt{a}}\] and see that you get \(a\)

OpenStudy (mww):

that last part is called rationalizing the denominator.

satellite73 (satellite73):

another is to use exponents \[\large \frac{a^1}{a^{\frac{1}{2}}}=a^{1-\frac{1}{2}}=a^{\frac{1}{2}}\]

OpenStudy (daisiablis11):

I see

satellite73 (satellite73):

and yet another way, as @mww said is to multiply top and bottom by \(\sqrt{a}\) aka rationalize the denominator

OpenStudy (daisiablis11):

I always get stuck with algebra. Thanks!

satellite73 (satellite73):

that is everyone's problem... yw